40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2354     Entry time: Tue Jun 4 20:44:21 2019
Author: rana 
Type: Summary 
Category: Other 
Subject: next steps 
  1. Fix the temperature control so that the fluctuation over hours time scale is low as Tara/Evan had it. This will allow for using the Marconi with a low FM range (or potentially none; we could just demod with a fixed frequency and directly digitize the IF signal in the CYMAC).
  2. Debug and fix the FSS misbehavior. Circuit creep in the last few years has degraded its performance relative to the in-loop performance achieved back in 2015.
  3. Recalculate the Brownian noise.
ELOG V3.1.3-